what is the solubility of CaF2 (Ksp = 1.46x10^-10) in a solution that has [F-] = 1.21M?
1
Expert's answer
2020-11-02T08:56:00-0500

Ksp=[Ca2+] * [F-]2

[Ca2+]=Ksp/[F-]2=1.46x10^(-10)/1.212=9.97*10^-11;

S=∛(Ksp/4)= 2.92 * 10^-4